Question 1.2 [4] An online survey of a large number of soccer fans determined that 65% follow European competitions and 55% follow local (South African) competitions while all of the fans follow at least one of the two activities. What proportion of soccer fans follow European competitions exclusively?


1
Expert's answer
2021-07-05T13:54:19-0400

A=follow European competitions

P(A)=0.65

B=follow local (American) competitions

P(B)=0.55

It is given that all of the fans follow at least one of the two activities. It means

"P(A \cup B) = 1"

By addition theorem

"P(A \cup B) = P(A) + P(B) -P(A \cap B) \\

P(A \cap B) = 0.65 + 0.55 -1 \\

P(A \cap B) = 0.2"

The proportion of soccer fans follow European competitions exclusively is computed as follows:

"= P(A) -P(A \cap B) \\

= 0.65-0.20 \\

= 0.45"

45% of soccer fans follow European competitions exclusively.